Maria bought seven boxes . A week later half of all her boxes were destroyed in a fire . There are only now 22 boxes left . With how many did she start with
Answer by ewatrrr(24785)

x - (x+7)/2 = 22 |Multiplying thru by 2so as all denominators = 1
2x - (x+7) = 44
x = 51, how many she started with
AND
51 - 58/2 = 51 - 29 = 22 CHECKS!